Find the following for f(x) = \frac{x^{2}+3x}{x−1} = \frac{x(x+3)}{x−1} (if they exist; if they don’t exist, state so). Use this information to graph f. Note f′(x) = \frac{(x−3)(x+1)}{(x−1)^{2}}, f′′(x) = \frac{8}{(x−1)^{3}} [1.5] 1a.) critical numbers: [1.5] 1b.) local maximum(s) occur at x = [1.5] 1c.) local minimum(s) occur at x = [1.5] 1d.) The global maximum of f on the interval [0, 5] is and occurs at x = [1.5] 1e.) The global minimum of f on the interval [0, 5] is and occurs at x = [1.5] 1f.) Inflection point(s) occur at x = [1.5] 1g.) f increasing on the intervals [1.5] 1h.) f decreasing on the intervals [1.5] 1i.) f is concave up on the intervals [1.5] 1j.) f is concave down on the intervals [1.5] 1k.) Equation(s) of vertical asymptote(s)

[4] 1l.) Equation(s) of horizontal and/or slant asymptote(s) [4.5] 1m.) Graph f